# Mixed Near-field and Far-field Target Localization for Low-altitude Economy

### v1

@@ III-D Extensions and Discussions
However, since the height of the peaks in $\mathbf{M}_{\rm DFT}$ depends on the signal powers $|g_{k}|$, when the received signal powers at the BS from far-field UAVs are relatively weak, their peaks in $\mathbf{M}_{\rm DFT}$ may be obscured by the side-lobes of near-field UAVs with stronger DFT-lobe signal power. In this case, the coarse estimation method based on 2D-DFT cannot be used to reduce the computational complexity of the proposed algorithm. How to overcome this issue will be left for our future work.
Figure 6: An example of normalized 2D-DFT of the constructed matrix $\mathbf{M}$. System parameters are the same as that in Fig. 5. Taking one peak index ($\hat{i}=62$, $\hat{j}=39$) as an example (4 peaks in total), the estimated parameter $\alpha$ are $\hat{\alpha}^{(1)}=\frac{62}{101}=0.6139$ and $\hat{\alpha}^{(2)}=\frac{62}{101}-1=-0.3861$, while parameter $\beta$ are estimated by $\hat{\beta}^{(1)}=\frac{39}{101}=0.3861$ and $\hat{\beta}^{(2)}=\frac{39}{101}-1=-0.6139$. The estimated angle parameters are corresponding to the near-field UAV located at ($-\frac{5\pi}{13}$, ${0.23\pi}$, $50$ m) where the parameter $\alpha=\sin(-\frac{5\pi}{13})\cos(0.23\pi)=-0.7014$ and $\beta=\sin(-\frac{5\pi}{13})\sin(0.23\pi)=-0.6183$. Obviously, the estimated parameter $\hat{\alpha}^{(1)}$ and $\hat{\beta}^{(1)}$ are corresponding ambiguous angles.
Passive sensing scenarios: The proposed mixed near-field and far-field target localization method can also be extended to the passive-sensing scenario where the BS estimates the UAV locations based on echo signals transmitted by itself [26]. For this case, let $\mathbf{x}_{\rm t}(m)=\mathbf{W}_{\rm t}\mathbf{s}(m)\in\mathbb{C}^{N\times 1}$ denote the transmitted signal by the BS at time $m$, where $\mathbf{W}_{\rm t}\in\mathbb{C}^{N\times N_{\rm RF}}$ represents the transmit combining matrix. Then, the echo signal (reflected by UAVs) received by the BS at discrete time $m$ is given by $\mathbf{y}(m)=\mathbf{W}_{\rm r}(m)\mathbf{G}\mathbf{x}(m)+\mathbf{W}_{\rm r}(m)\mathbf{n}(m)$, where $\mathbf{W}_{\rm r}(m)$ represents receive combining matrix at time $m$ and $\mathbf{x}(m)=[x_{1}(m),x_{2}(m),\cdots,x_{K}(m)]^{T}\in\mathbb{C}^{K\times 1}$ denotes the equivalent transmitted signal vector of the $K$ UAVs. In particular, $x_{k}(m)\in\mathbb{C}$ is the equivalent transmitted signal at $k$-th UAV, given by
$x_{k}(m)=\gamma_{k}\mathbf{b}^{H}(r_{k},\theta_{k},\phi_{k})\mathbf{W}_{\rm t}\mathbf{s}(m),\leavevmode\nobreak\ \leavevmode\nobreak\ \forall k\in\mathcal{K},$ (47)
@@
It is worth noting that when UAVs are widely separated and $N_{\rm RF}$ is large enough, the correlation between $\widetilde{\mathbf{b}}^{H}(r_{i},\theta_{i},\phi_{i})$ and $\widetilde{\mathbf{b}}^{H}(r_{j},\theta_{j},\phi_{j})$ ($i\neq j$) is negligible. As a result, $\mathbf{R}_{x}$ is approximately a diagonal matrix, i.e., $\mathbf{R}_{x}\approx\mathrm{diag}\{|\gamma_{1}|^{2}N_{\rm RF},\cdots,|\gamma_{k}|^{2}N_{\rm RF},\cdots,|\gamma_{K}|^{2}N_{\rm RF}\}$. Hence, the original covariance matrix of the echo signals in (48) has the same form as the covariance matrix of passive signals in (18), for which our proposed mixed-field target localization method can be directly applied. However, when channel correlation between different UAVs is large enough, the anti-diagonal elements of $\mathbf{R}$ will be a linear combination of the signals received by multiple antennas, leading to erroneous peaks in the corresponding angular spectrum. How to address this issue will be left for our future work.
@@ IV Cramér-Rao bound derivation
In this section, we characterize the CRB for the mixed-field parameter estimation for evaluating the performance of the proposed UAV localization algorithm.
@@
$\mathrm{MSE}([\hat{\bm{\eta}}]_{i})={\mathbb{E}}\Big{\{}\left([\hat{\bm{\eta}}]_{i}-[\bm{\eta}]_{i}\right)^{2}\Big{\}}\geq\left[\mathbf{CRB}(\bm{\eta})\right]_{i,i},$ (50)
where $\bm{\eta}$ represents the parameter vector and $\mathbf{CRB}(\bm{\eta})$ denotes the corresponding CRB matrix, which can be obtained by the inversion of the Fisher information matrix (FIM). In this paper, the parameter vector of interest is given by
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \phi_{K_{1}+1},\cdots,\phi_{K},r_{K_{1}+1},\cdots,r_{K}]^{T}.$
@@
As such, we obtain the matrix closed-form CRB for the mixed-field localization. In particular, when all the UAVs are located in the far or near-field, the CRB matrices reduces to $\mathbf{CRB}(\bm{\eta})=\mathbf{F}_{\rm FF}^{-1}$ and $\mathbf{CRB}(\bm{\eta})=\mathbf{F}_{\rm NN}^{-1}$, respectively.
@@ V-A System Setup and Benchmark Schemes
System parameters are set as follows, unless otherwise specified. The BS operates at the $f=10$ GHz frequency band, and the UPA at the BS is equipped with $61\times 61$ antennas. In addition, the number of RF chains is $N=61$ with $M_{x}=61$ and $M_{y}=1$. The number of equivalent snapshots is set as $L=500$. There are four targets in the setup: 1) far-field UAV 1 located at $(\frac{\pi}{4},-\frac{\pi}{3})$; 2) far-field UAV 2 located at $(\frac{\pi}{8},\frac{\pi}{3})$; near-field UAV 3 located at ($\frac{\pi}{4}$,$\frac{\pi}{4}$, $30$ m); near-field UAV 4 located at ($\frac{\pi}{8}$,$\frac{\pi}{4}$, $40$ m). Moreover, we consider the RMSE for angle and range estimation for performance evaluation, which are defined by $\theta_{\mathrm{RMSE}}=\Big{(}{\frac{1}{QK}\sum_{q=1}^{Q}\sum_{k=1}^{K}\big{(}\hat{\theta}_{k}^{(q)}-\theta_{k}\big{)}^{2}}\Big{)}^{\frac{1}{2}}$, $\phi_{\mathrm{RMSE}}=\Big{(}{\frac{1}{QK}\sum_{q=1}^{Q}\sum_{k=1}^{K}\big{(}\hat{\phi}_{k}^{(q)}-\phi_{k}\big{)}^{2}}\Big{)}^{\frac{1}{2}}$ and $r_{\mathrm{RMSE}}=\Big{(}{\frac{1}{QK}\sum_{q=1}^{Q}\sum_{k=1}^{K}\big{(}\hat{r}_{k}^{(q)}-r_{k}\big{)}^{2}}\Big{)}^{\frac{1}{2}}$, where $\hat{\theta}_{k}^{(q)}$, $\hat{\phi}_{k}^{(q)}$ and $\hat{r}_{k}^{(q)}$ denote the estimated elevation and azimuth angles and range for the $k$-th UAV in the $q$-th Monte Carlo simulation over $Q=100$ Monte Carlo simulations. For performance comparison, we consider the following benchmark schemes:
@@
  • Proposed method with ${d_{0}=\frac{\lambda}{4}}$: This scheme is similar to our proposed method. It does has the angular ambiguity issue, but it has smaller array aperture and hence potentially reduced localization accuracy.
@@
Table I: UAV classification under different methods.
Method UAV index Far-field UAV 1 $(\frac{\pi}{4},\frac{\pi}{3},-)$ Far-field UAV 2 $(\frac{\pi}{8},\frac{\pi}{3},-)$ Near-field UAV 3 ($\frac{\pi}{4}$,$\frac{\pi}{4}$, $30$ m) Near-field UAV 4 ($\frac{\pi}{8}$,$\frac{\pi}{4}$, $40$ m)
@@ V-B Performance Analysis
In Table I, we summarize the effectiveness of different mixed-field classification methods, where the reference SNR is set as 10 dB. It is observed that the proposed method and the benchmark scheme in [17] achieve similar accuracy in angle estimation. However, the method proposed in [17] wrongly estimate the near-field UAVs as far-field UAVs and thus cannot estimate the ranges of near-field UAVs (see Section III-D).
@@
Finally, we plot the running time versus the number of antennas in Fig. 11, where the numbers of search grids are set as $\ell_{x}=\ell_{y}=10^{4}$ and $\ell_{d}=10^{3}$. It is observed that the proposed method as well as the 2D-DFT method significantly reduce the computational complexity. This is because the proposed method decouple the angle and range estimation, while the 2D-DFT method first coarsely estimates UAV angles and then refine angle estimation by applying 1D MUSIC algorithm, which narrows the searching ranges.
@@ VI Conclusions
In this paper, we proposed a three-step mixed near-field and far-field target localization method by employing typical wireless communication infrastructures. First, we judiciously designed the analog combining matrix at the BS over time to recover the received signals at antennas with minimum recovery errors. Then, we decoupled the angle and range estimation problems via extracting the anti-anti-diagonal elements of the original covariance matrix, for which a decoupled 2D MUSIC algorithm was proposed to estimate UAV angles including the true far-field and near-field angles as well as their corresponding ambiguous angles. Further, an effective classification method was proposed to distinguish mixed-field targets and resolve the angular ambiguity based on different patterns in the range-domain MUSIC spectrum of the three types of angles. Finally, we derived the CRB for mixed-field target localization as the estimation error lower bound and numerical results were presented to demonstrate the effectiveness of the proposed method as compared with various benchmark schemes.

### v2

@@ III-D Extensions and Discussions
However, since the height of the peaks in $\mathbf{M}_{\rm DFT}$ depends on the signal powers $|g_{k}|$, when the received signal powers at the BS from far-field targets are relatively weak, their peaks in $\mathbf{M}_{\rm DFT}$ may be obscured by the side-lobes of near-field targets with stronger DFT-lobe signal power. In this case, the coarse estimation method based on 2D-DFT cannot be used to reduce the computational complexity of the proposed algorithm. How to overcome this issue will be left for our future work.
Figure 6: An example of normalized 2D-DFT of the constructed matrix $\mathbf{M}$. System parameters are the same as that in Fig. 5. Taking one peak index ($\hat{i}=62$, $\hat{j}=39$) as an example (4 peaks in total), the estimated parameter $\alpha$ are $\hat{\alpha}^{(1)}=\frac{62}{101}=0.6139$ and $\hat{\alpha}^{(2)}=\frac{62}{101}-1=-0.3861$, while parameter $\beta$ are estimated by $\hat{\beta}^{(1)}=\frac{39}{101}=0.3861$ and $\hat{\beta}^{(2)}=\frac{39}{101}-1=-0.6139$. The estimated angle parameters are corresponding to the near-field target located at ($\frac{5\pi}{13}$, $0.23\pi$}, $50$ m) where the parameter $\alpha=\sin(-\frac{5\pi}{13})\cos(0.23\pi)=-0.7014$ and $\beta=\sin(-\frac{5\pi}{13})\sin(0.23\pi)=-0.6183$. Obviously, the estimated parameter $\hat{\alpha}^{(1)}$ and $\hat{\beta}^{(1)}$ are corresponding ambiguous angles.
Active sensing scenarios: The proposed mixed near-field and far-field target localization method can also be extended to the active-sensing scenario where the BS estimates the target locations based on echo signals transmitted by itself [19]. For this case, let $\mathbf{x}_{\rm t}(m)=\mathbf{W}_{\rm t}\mathbf{s}(m)\in\mathbb{C}^{N\times 1}$ denote the transmitted signal by the BS at time $m$, where $\mathbf{W}_{\rm t}\in\mathbb{C}^{N\times N_{\rm RF}}$ represents the transmit combining matrix. Then, the echo signal (reflected by targets) received by the BS at discrete time $m$ is given by $\mathbf{y}(m)=\mathbf{W}_{\rm r}(m)\mathbf{G}\mathbf{x}(m)+\mathbf{W}_{\rm r}(m)\mathbf{n}(m)$, where $\mathbf{W}_{\rm r}(m)$ represents receive combining matrix at time $m$ and $\mathbf{x}(m)=[x_{1}(m),x_{2}(m),\cdots,x_{K}(m)]^{T}\in\mathbb{C}^{K\times 1}$ denotes the equivalent transmitted signal vector of the $K$ targets. In particular, $x_{k}(m)\in\mathbb{C}$ is the equivalent transmitted signal at $k$-th target, given by
$x_{k}(m)=\gamma_{k}\mathbf{b}^{H}(r_{k},\theta_{k},\phi_{k})\mathbf{W}_{\rm t}\mathbf{s}(m),\penalty 10000\ \penalty 10000\ \forall k\in\mathcal{K},$ (47)
@@
It is worth noting that when targets are widely separated and $N_{\rm RF}$ is large enough, the correlation between $\widetilde{\mathbf{b}}^{H}(r_{i},\theta_{i},\phi_{i})$ and $\widetilde{\mathbf{b}}^{H}(r_{j},\theta_{j},\phi_{j})$ ($i\neq j$) is negligible. As a result, $\mathbf{R}_{x}$ is approximately a diagonal matrix, i.e., $\mathbf{R}_{x}\approx\mathrm{diag}\{|\gamma_{1}|^{2}N_{\rm RF},\cdots,|\gamma_{k}|^{2}N_{\rm RF},\cdots,|\gamma_{K}|^{2}N_{\rm RF}\}$. Hence, the original covariance matrix of the echo signals in (48) has the same form as the covariance matrix of signals in (18), for which our proposed mixed-field target localization method can be directly applied. However, when channel correlation between different targets is large enough, the anti-diagonal elements of $\mathbf{R}$ will be a linear combination of the signals received by multiple antennas, leading to erroneous peaks in the corresponding angular spectrum. How to address this issue will be left for our future work.
@@ IV Cramér-Rao bound derivation
In this section, we characterize the CRB for the mixed-field parameter estimation for evaluating the performance of the proposed localization algorithm.
@@
$\mathrm{MSE}([\hat{\bm{\eta}}]_{i})\!=\!{\mathbb{E}}\Big\{\left([\hat{\bm{\eta}}]_{i}\!-\![\bm{\eta}]_{i}\right)^{2}\Big\}\!\geq\!\left[\mathbf{CRB}(\bm{\eta})\right]_{i,i},\forall i\in\mathcal{I},$
where $\bm{\eta}$ represents the parameter vector and $\mathbf{CRB}(\bm{\eta})$ denotes the corresponding CRB matrix, which can be obtained by the inversion of the Fisher information matrix (FIM). In addition, $\mathcal{I}\triangleq\{1,2,\cdots,2K+K_{2}\}$ denotes the index set for parameters to be estimated. In this paper, the parameter vector of interest is given by
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \phi_{K_{1}+1},\cdots,\phi_{K},r_{K_{1}+1},\cdots,r_{K}]^{T}.$
@@
As such, we obtain the matrix closed-form CRB for the mixed-field localization. In particular, when all the targets are located in the far or near-field, the CRB matrices reduces to $\mathbf{CRB}(\bm{\eta})=\mathbf{F}_{\rm FF}^{-1}$ and $\mathbf{CRB}(\bm{\eta})=\mathbf{F}_{\rm NN}^{-1}$, respectively.
@@ V-A System Setup and Benchmark Schemes
System parameters are set as follows, unless otherwise specified. The BS operates at the $f=10$ GHz frequency band, and the UPA at the BS is equipped with $61\times 61$ antennas, for which the Rayleigh distance can be calculated as $Z_{\rm R}=2\times(61^{2}\times 0.015^{2}+61^{2}\times 0.015^{2})/{0.03}\approx 112$ m. In addition, the number of RF chains is $N=61$ with $M_{x}=61$ and $M_{y}=1$. The number of equivalent snapshots is set as $L=500$. There are four targets in the setup: 1) far-field target 1 located at $(\frac{\pi}{4},-\frac{\pi}{3})$ rad; 2) far-field target 2 located at $(\frac{\pi}{8},\frac{\pi}{3})$ rad; near-field target 3 located at ($\frac{\pi}{4}$ $\frac{\pi}{4} rad$, $30$ m); near-field target 4 located at ($\frac{\pi}{8}$ $\frac{\pi}{4} rad$, $40$ m). Moreover, we consider the RMSE for angle and range estimation for performance evaluation, which are defined as $\theta_{\mathrm{RMSE}}=\Big({\frac{1}{QK}\sum_{q=1}^{Q}\sum_{k=1}^{K}\big(\hat{\theta}_{k}^{(q)}-\theta_{k}\big)^{2}}\Big)^{\frac{1}{2}}$, $\phi_{\mathrm{RMSE}}=\Big({\frac{1}{QK}\sum_{q=1}^{Q}\sum_{k=1}^{K}\big(\hat{\phi}_{k}^{(q)}-\phi_{k}\big)^{2}}\Big)^{\frac{1}{2}}$ and $r_{\mathrm{RMSE}}=\Big({\frac{1}{QK}\sum_{q=1}^{Q}\sum_{k=1}^{K}\big(\hat{r}_{k}^{(q)}-r_{k}\big)^{2}}\Big)^{\frac{1}{2}}$, where $\hat{\theta}_{k}^{(q)}$, $\hat{\phi}_{k}^{(q)}$ and $\hat{r}_{k}^{(q)}$ denote the estimated elevation and azimuth angles and range for the $k$-th target in the $q$-th Monte Carlo simulation over $Q=100$ Monte Carlo simulations. For performance comparison, we consider the following benchmark schemes:
@@
  • Localization method $d_{0}=\frac{\lambda}{4}$}: This scheme is similar to our proposed method. It does has the angular ambiguity issue, but it has smaller array aperture and hence potentially reduced localization accuracy.
@@
Table I: Target classification under different methods.
Method Target index Far-field target 1 $(\frac{\pi}{4},\frac{\pi}{3},-)$ Far-field target 2 $(\frac{\pi}{8},\frac{\pi}{3},-)$ Near-field target 3 ($\frac{\pi}{4}$$\frac{\pi}{4}$, $30$ m) Near-field target 4 ($\frac{\pi}{8}$$\frac{\pi}{4}$, $40$ m)
@@ V-B Performance Analysis
In Table I, we summarize the effectiveness of different mixed-field classification methods, where the reference SNR of each target (defined by $\frac{N|\beta_{k}|^{2}p_{k}}{r_{k}^{2}\sigma^{2}}$) is set as 10 dB. It is observed that the proposed method and the benchmark scheme in [6] achieve similar accuracy in angle estimation. However, the method proposed in [6] wrongly estimate the near-field targets as far-field targets and thus cannot estimate the ranges of near-field targets (see Section III-D).
@@
Finally, we plot the running time versus the number of antennas in Fig. 11, where the numbers of search grids are set as $\ell_{x}=\ell_{y}=10^{4}$ and $\ell_{d}=10^{3}$. It is observed that the proposed method as well as the “proposed method + DFT” significantly reduce the computational complexity. This is because the proposed method decouple the angle and range estimation, while the 2D-DFT method first coarsely estimates target angles and then refine angle estimation by applying 1D MUSIC algorithm, which narrows the search regime.
@@ VI Conclusions
In this paper, we proposed a three-step mixed near-field and far-field target localization method by employing typical wireless communication infrastructures. First, we judiciously designed the analog combining matrix at the BS over time to recover the received signals at antennas with minimum recovery errors. Then, we decoupled the angle and range estimation problems via extracting the anti-anti-diagonal elements of the original covariance matrix, for which a decoupled 2D MUSIC algorithm was proposed to estimate target angles including the true far-field and near-field angles as well as their corresponding ambiguous angles. Further, an effective classification method was proposed to distinguish mixed-field targets and resolve the angular ambiguity based on different patterns in the range-domain MUSIC spectrum of the three types of angles. Finally, we derived the CRB for mixed-field target localization as the estimation error lower bound and numerical results were presented to demonstrate the effectiveness of the proposed method as compared with various benchmark schemes.